OpenStudy (anonymous):

The gas tank of a truck is a cylinder 5ft long with a diameter of 2ft. At the gas station, a pump pours gas at the rate of 4ft^3 per minute. How many minutes does it take to fill the empty tank with that pump?

OpenStudy (wolf1728):

Volume of a cylinder = PI*r²*h Volume = PI*1²*5 = 15.71 cubic feet If gas pump fills at 4 cubic ft / minute then it takes 15.71 / 4 = 3.9275 minutes about 4 minutes